OpenStudy (anonymous):

what order pair is NOT a solution of the equation y=x-3? f. (-4,-7) g. (12,9) h.(0,-3) i. (-8,11)

OpenStudy (anonymous):

Just substitute each combo and check.

OpenStudy (anonymous):

im not getting it @douglaswinslowcooper

OpenStudy (anonymous):

y = x - 3 Try x = -4, y = -7 Check: y = (-4) - 3 = -7 so this pair works Does I work?

OpenStudy (anonymous):

yes @douglaswinslowcooper

OpenStudy (anonymous):

is it g ? @douglaswinslowcooper